Estimate the distance traveled when driving 48 miles per hour for 7 hours, 8 minutes

Answer by mananth(16949)

7 hours 8 minutes
8 minutes = 8/60 = 2/15 hours
..
7+2/15
=107/15 hours
..
d= rt
d= 48 * 107/15
d= 342.4 miles